### v3

# Explicit surjectivity of Galois representations of products of elliptic curves over function fields

Abstract.
We prove an explicit surjectivity result for products of non-isotrivial, non-isogenous elliptic curves over a function field of arbitrary characteristic. This is by way of an isogeny degree bound in this setting, generated from bounds for elliptic curves by Griffon–Pazuki, and techniques originated by Serre and Masser–Wüstholz in the number field setting. We apply our result to prove that most members of a family of products of elliptic curves over $\mathbb{Q}$ with no extra endomorphisms have no exceptional primes above a specified constant which depends neither on the elliptic curve factors nor on the dimension of the product.
Key words and phrases:
Galois representation, elliptic curve, function field, abelian variety, isogeny
2020 Mathematics Subject Classification:
Primary 11G10, Secondary 11G05
1. Introduction
Let $E$ be an elliptic curve over a field $K$, and let $N$ be a positive integer. The absolute Galois group $G_{K}:=\operatorname{Gal}(K^{\textnormal{sep}}/K)$ acts naturally on the torsion points of $E$ over a separable closure $K^{\textnormal{sep}}$ of $K$, and in particular acts on the subgroup $E[N]$ of points of order dividing $N$. This action is captured in the mod-$N$ Galois representation
$\rho_{E,N}:G_{K}\rightarrow\operatorname{GL}(E[N])\simeq\operatorname{GL}_{2}(\mathbb{Z}/N\mathbb{Z}).$
By a celebrated result of Serre [Ser72, Théorème 2], if $K$ is a number field and $E$ does not have complex multiplication over $K^{\textnormal{sep}}$ (is non-CM), then there exists a constant $\ell_{0}$ depending on $E$ such that $\rho_{E,\ell}$ is surjective for all primes $\ell\geq\ell_{0}$. Serre also proved an analogous result [Ser72, Théorème 6] for a product of two elliptic curves, stating that the mod-$\ell$ Galois representation associated to a product of two non-CM, non-(geometrically)-isogenous elliptic curves over $K$ is as large as possible for all $\ell$ sufficiently large with respect to the two elliptic curve factors.
For $K$ the function field of a smooth, projective, and geometrically integral curve over a perfect field of constants $k$, we have a theorem of Igusa [Igu59, Theorem 3] whose proof preceded Serre’s theorem. (We also refer the reader to [BLV09] for an alternate proof of this result.) Igusa’s theorem implies that if $E$ is a non-isotrivial elliptic curve over $K$, then there exists a constant $\ell_{0}$ depending on $E$ such that for all $\ell\geq\ell_{0}$ the representation $\rho_{E,\ell}$ is as large as possible in the sense that
$\text{Image}(\rho_{E,\ell})\cap\operatorname{SL}_{2}(\mathbb{Z}/N\mathbb{Z})=\operatorname{SL}_{2}(\mathbb{Z}/N\mathbb{Z}).$
Naturally, one asks for effective versions of the above results: what explicit lower bounds on $\ell$ guarantee that the image is as large as possible in either setting? Following Serre’s work, an effective theorem for the product of an arbitrary number of elliptic curves over a number field was proven by Masser–Wüstholz [MW93a, Theorem, Proposition 1] as an application of their work on isogeny degree bounds in [MW93b]. They prove that surjectivity holds for all primes $\ell$ larger than a constant depending on the Weil heights of the elliptic curve factors and the degree of the number field. Since their work, the subject of effective surjectivity results over number fields has seen ongoing interest, with many working to improve the bounds of Masser–Wüstholz or to extend these types of results to other classes of abelian varieties (see, e.g., [Coj05, LV14, Lom15, LF16, Lom16b, Lom16a, Zyw22, MW24b, MW24a]).
An explicit, effective surjectivity result for elliptic curves over a function field, which we recall as Theorem 4.1, was proven by Cojocaru–Hall in [CH05]. Their theorem states that surjectivity holds for all primes $\ell>c(g)$ not equal to the characteristic of $K$, where $c(g)$ is an explicit constant depending only on the genus of $K$. (The precise definition of $c(g)$ appears both in Theorem 1.1 and in Theorem 4.1.)
Our main result in this work is an explicit surjectivity result for the product of an arbitrary number of elliptic curves over a function field, which provides the first extension of both Igusa’s theorem and Cojocaru–Hall’s theorem to products of dimension larger than $1$. For $C$ a smooth, projective, and geometrically integral curve over a perfect field $k$, let $K=k(C)$ denote the function field of $C$ and let $g=g(K)=g(C)$ denote the genus of $K$. For $E_{1},\ldots,E_{n}$ elliptic curves over $K$,
$A=E_{1}\times\cdots\times E_{n}$
the product abelian variety of dimension $n$, and $\ell$ a prime, let
$\rho_{A,\ell}:G_{K}$ $\rightarrow\operatorname{GL}(E_{1}[\ell]\times\cdots\times E_{n}[\ell])\simeq\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}$
be the mod-$\ell$ Galois representation of $A$. Equivalently, this is the product of the mod-$\ell$ Galois representations of the factors $E_{i}$. The determinants of the projections of $\rho_{A,\ell}$ to each component must agree, each being equal to the $\ell^{\textnormal{th}}$ cyclotomic character of $K$, and so the image of $\rho_{A,\ell}$ is contained in the determinant locus
$\Delta_{\ell}:=\left\{(M_{1},M_{2},\ldots,M_{n})\mid\det{M_{1}}=\det{M_{2}}=\ldots=\det{M_{n}}\right\}\leq\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}.$
Let $\Psi_{\ell}:=\textnormal{Image}(\rho_{A,\ell})\cap\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}$. Our effective result for products is then as follows (see Theorem 4.3).
Theorem 1.1.
Let $K$ be the function field of a smooth, projective, and geometrically integral curve $C$ of genus $g$ over a perfect field $k$. Let $E_{1},\ldots,E_{n}$ be pairwise non-isogenous elliptic curves over $K$ with $j(E_{i})\not\in\overline{k}$ for each $1\leq i\leq n$, and let
$A=E_{1}\times\cdots\times E_{n}.$
Set
$e_{2}(q):=\begin{cases}1\quad&\text{ if }q\equiv 1\pmod{4}\\ -1\quad&\text{ otherwise,}\end{cases}\quad\quad e_{3}(q):=\begin{cases}1\quad&\text{ if }q\equiv 1\pmod{3}\\ -1\quad&\text{ otherwise,}\end{cases}\quad$
$c(g)$ $:=2+\max\left\{q\textnormal{ prime }\bigm|\dfrac{q-(6+3e_{2}(q)+4e_{3}(q))}{12}\leq g\right\},$
$C(g)$ $=3176523\cdot\max\{1,g^{3/2}\},$
and finally
$\widetilde{C}(g):=\begin{cases}c(g)\quad\text{ if }n=1\\ C(g)\quad\text{ if }n\geq 2.\end{cases}$
For all primes $\ell$ not equal to $\textnormal{char}(K)$ satisfying $\ell>\widetilde{C}(g)$, we have $\Psi_{\ell}=\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}$.
A pleasing aspect of Theorem 1.1, like the dimension $1$ result Theorem 4.1, is that it depends only on the genus $g$ of $K$, and not on the individual elliptic curves factors nor on the dimension of $A$ when $n\geq 2$. The constant $\widetilde{C}$ appearing in the above statement is defined in cases, singling out the dimension $1$ case, so as to recover the theorem of Cojocaru–Hall when $n=1$, given that our proof strategy only produces a bound in dimension $n\geq 2$ (see Theorem 4.3). We also prove a version (Theorem 4.5) of this result which depends on the modular heights of the factors, which can provide a better bound than $C(g)$ for individual product abelian varieties of dimension $n\geq 2$.
We apply Theorem 1.1 to prove a result on exceptional primes of specializations of families of products of elliptic curves over the rationals. Our setup is as follows: Let $n\geq 2$, let $A_{1}(t),B_{1}(t),\ldots,A_{n}(t),$ and $B_{n}(t)\in\mathbb{Z}[t]$ and consider the elliptic curves over $\mathbb{Q}(t)$ given by
$E_{i}:y^{2}=x^{3}+A_{i}(t)x+B_{i}(t)\qquad\textnormal{ for i }\in\{1,\ldots,n\}.$
Assume that the $E_{i}$ are pairwise non-isogenous over $\overline{\mathbb{Q}(t)}$ and that $j(E_{i})\not\in\overline{\mathbb{Q}}$ for each $i$. Let
$A=E_{1}\times\cdots\times E_{n},$
and let $E_{i,t_{0}}$ and $A_{t_{0}}$ denote, respectively, the specialization of $E_{i}$ and of $A$ to a given $t_{0}\in\mathbb{Q}$. Let $S\subseteq\mathbb{Q}$ denote the set of $t_{0}\in\mathbb{Q}$ such that at least one of the following items holds:
  • AA has bad reduction at $t_{0}$,
  • $E_{i,t_{0}}$ has CM for some $i\in\{1,\ldots,n\}$, or
  • $E_{i,t_{0}}$ and $E_{j,t_{0}}$ are isogenous over $\overline{\mathbb{Q}}$ for some $i\neq j$.
For $T>0$ a real number, let
$\mathcal{F}(T):=\left\{\frac{m}{n}\in\mathbb{Q}\setminus S\mid m,n\in\mathbb{Z},\gcd(m,n)=1,\textnormal{ and }\max\{|m|,|n|\}\leq T\right\},$
and for $\ell$ a prime number let
$\mathcal{E}_{\ell}(T):=\{t_{0}\in\mathcal{F}(T)\mid\operatorname{Gal}(\mathbb{Q}(A_{t_{0}}[\ell])/\mathbb{Q})\subsetneq\Delta_{\ell}\}.$
Our result, stating that most specializations of $A$ have no exceptional primes at least $3176533$, is as follows (see Theorem 5.5).
Theorem 1.2.
Letting
$\mathcal{E}(T):=\bigcup_{\ell\geq 3176533}\mathcal{E}_{\ell}(T),$
we have
$\lim_{T\to\infty}\dfrac{|\mathcal{E}(T)|}{|\mathcal{F}(T)|}=0.$
Our proof of Theorem 1.1 follows the techniques of Masser–Wüstholz from [MW93a], which in our setting require as input an isogeny degree estimate for product abelian surfaces over function fields. This comes in the form of Proposition 3.4, which we prove by bootstrapping isogeny degree estimates for elliptic curves over function fields obtained in recent work of Griffon–Pazuki [GP22, Proposition 6.5, Proposition 6.7]. Our argument, like that of Griffon–Pazuki in the dimension $1$ case, involves points on modular curves $X_{0}(N)$, of specified level $N$, induced by minimal-degree isogenies along with genus bounds for $X_{0}(N)$ in terms of the level $N$. Our proof of Theorem 1.2 is along the same lines as the proof of Cojocaru–Hall’s result on families of elliptic curves over $\mathbb{Q}$ [CH05, Theorem 1.3], making use of Gallagher’s large sieve.
The paper is structured as follows. In Section 2, we briefly review necessary preliminaries related to function fields, heights, and isogenies. We prove isogeny degree bounds in Section 3, and apply one such bound to prove Theorem 1.1 in Section 4. We prove results on families of products, in particular Theorem 1.2, in Section 5.
Acknowledgments
We thank Oana Padurariu and Nathan Jones for feedback on earlier drafts of this work, and we thank Ariel Weiss for a conversation that helped us to improve our main result Theorem 1.1; see Remark 4.4.
2. Background material
2.1. Function fields and heights
Let $C$ be a smooth, projective, and geometrically integral curve of genus $g$ over a perfect field $k$ of characteristic $p=\textnormal{char}(K)\geq 0$. We will let $K=K(C)$ denote the function field of $C$, and set $g=g(K):=g(C)$ to be the genus of $K$ (which is independent of the choice of model $C$). In this work, by a function field we always mean $K$ of the above form with field of constants $k$. We refer the reader to [Ros02] for general background on function fields.
Let $L/K$ be a finite degree extension. For each place $w$ of $L$, let $|\cdot|_{w}$ denote the corresponding normalized absolute value on $L$. Let $v$ denote the unique place of $K$ lying below $w$ and let $L_{w}$ and $K_{v}$, respectively, denote the corresponding completions of $L$ and $K$. The local degree of $w$ is the quantity
$n_{w}:=[L_{w}:k_{v}].$
Let $P=[x_{0}:x_{1}]\in\mathbb{P}^{1}(L)$. The height of $P$ with respect to $L$ is
$h_{L}(P):=\sum_{w\textnormal{ a place of }L}n_{w}\log\max\{|x_{0}|_{w},|x_{1}|_{w}\}.$
We now define the Weil height of $P$, which is independent of the choice of extension $L$ such that $P\in\mathbb{P}^{1}(L)$.
Definition 2.1.
The (absolute logarithmic) Weil height of $P\in\mathbb{P}^{1}(L)$ is
$h(P):=\frac{h_{L}(P)}{[L:K]}.$
For $f\in\overline{K}$, the Weil height of $f$ is then
$h(f):=h([f:1]).$
Like results of [MW93a] and [GP22], versions of our effective results on products and isogeny degree bounds will involve the modular heights of the elliptic curve factors. We recall the definition from [GP22, §1].
Definition 2.2.
For $E$ an elliptic curve over $\overline{K}$ with $j$-invariant $j(E)$, the modular height of $E$ is the (absolute logarithmic) Weil height of $j(E)$:
$\mathrm{h}_{\textnormal{mod}}(E):=h(j(E)).$
2.2. Isogenies
For general background on abelian varieties, we recommend [EvdGM, Mil86, Mum08]. We will most often reference [EvdGM], which from the start works over a general (not necessarily algebraically closed) base field when possible.
Unless otherwise specified, in this work isogenies will be assumed to be defined over $\overline{K}$ (equivalently, over some finite extension of $K$).
For abelian varieties $X/K$ and $Y/K$ of the same dimension with respective identity elements $0_{X}$ and $0_{Y}$, we let $\textnormal{Hom}_{\overline{K}}(X,Y)$ denote the ring consisting of all isogenies $\left(X\otimes_{\operatorname{Spec}K}\operatorname{Spec}\overline{K}\right)\rightarrow Y\otimes_{\operatorname{Spec}K}\operatorname{Spec}\overline{K}$ between $X$ and $Y$ over $\overline{K}$ along with the constant map sending every element of $X$ to $0_{Y}$. For $X/K$ a fixed abelian variety, we let $\textnormal{End}_{\overline{K}}(X)=\textnormal{Hom}_{\overline{K}}(X,X)$. The degree $\textnormal{deg}(\varphi)$ of an isogeny $\varphi:X\rightarrow Y$ is the degree $K(X):\varphi^{*}(K(Y))]$ of the corresponding extension of function fields. Equivalently, $\textnormal{deg}(\varphi)$ is the rank of the finite group scheme $\textnormal{ker}(\varphi)$ [EvdGM, p. 73].
The following result is due to Deuring in positive characteristic, and its statement and a proof which is agnostic to the characteristic can be found as [GP22, Lemma 2.1].
Lemma 2.3.
Let $K$ be a function field and $E/K$ an elliptic curve with $j(E)\not\in\overline{k}$. Then $\textnormal{End}_{\overline{K}}(E)\simeq\mathbb{Z}$.
We call an elliptic curve $E$ over a function field $K$ isotrivial if $j(E)\in\overline{k}$, and we call $E$ non-isotrivial otherwise. An elliptic curve with CM over a number field must have integral $j$-invariant, and the above lemma should be read as an analogue over functions fields: if $E/K$ is non-isotrivial (has non-constant $j$-invariant), then it does not have extra endomorphisms.
2.2.1. Factorizations
The main thrust of this subsection is to define, for an isogeny $\varphi$, an auxiliary isogeny in the other direction which will be relevant to our definition of biseparability to come. Some of the work here is not strictly necessary for the remainder of this work, but is useful in identifying how one might seek to generalize our isogeny-degree bound in Section 3.
Proposition 2.4.
Let $X,Y$ be abelian varities over $K$, and let $\varphi\in\textnormal{Hom}_{\overline{K}}(X,Y)$. Let $H\subseteq\textnormal{ker}(\varphi)$ be a subgroup scheme, and let $q:X\rightarrow X/H$ denote the natural quotient map. There exists an isogeny $\psi:X/H\rightarrow Y$ such that we get a factorization of isogenies $\varphi=\psi\circ q$.
Proof.
The main claim here is the implicit one that $X/H$ is again an abelian variety, which follows from [EvdGM, (4.39) Theorem]. We then get the isogeny $\psi$ from the universal property for quotients. ∎
The following result will be useful when we need certain factorizations of isogenies to be unique.
Lemma 2.5.
[EvdGM, (5.4) Lemma] Let $f:W\rightarrow X$ and $h:Y\rightarrow Z$ be isogenies of abelian varieties over a field $k$. If $g_{1},g_{2}:X\rightarrow Y$ are isogenies such that $h\circ g_{1}\circ f=h\circ g_{2}\circ f$, then $g_{1}=g_{2}$.
We now introduce, for an isogeny $\varphi$, an associated isogeny in the opposite direction which we label $\widetilde{\varphi}$.
Proposition 2.6.
If $\varphi:X\rightarrow Y$ is an isogeny of degree $d$, then there exists a unique isogeny $\widetilde{\varphi}:Y\rightarrow X$ such that
$\widetilde{\varphi}\circ\varphi=[d]_{X}\qquad\textnormal{ and }\qquad\varphi\circ\widetilde{\varphi}=[d]_{Y}.$
Proof.
This is [EvdGM, (5.12) Proposition]. The uniqueness claim is not explicitly made there, but it follows immediately from Lemma 2.5
When the dimension of $X$ and $Y$ is greater than $1$, the isogeny $\widetilde{\varphi}$ is not the dual isogeny of $\varphi$ (note that we have taken care to say nothing about polarizations). In particular, one can see by comparing degrees that the degree of $\varphi$ will be less than that of $\widetilde{\varphi}$ in dimension greater than $1$, and as a result $\widetilde{\widetilde{\varphi}}:X\rightarrow Y$ will not be equal to $\varphi$. More specifically, we note the following properties. These can be found in [Ros86] for isogenies over $\mathbb{C}$. Note that Rosen does refer to $\widetilde{\varphi}$ as the “dual isogeny,” whereas we would label this a misnomer.
Proposition 2.7.
Let $\varphi:X\rightarrow Y$ be an isogeny of degree $d$ between abelian varieties of dimension $n$, and let $\varphi^{\prime}:Y\rightarrow Z$ be an isogeny of degree $d^{\prime}$. We then have the following properties:
  1. (1)
    $\textnormal{deg}\left(\widetilde{\varphi}\right)=d^{2n-1}$.
  2. (2)
    $\widetilde{\widetilde{\varphi}}=[d^{2n-2}]_{Y}\circ\varphi$.
  3. (3)
    $\widetilde{\varphi^{\prime}\circ\varphi}=\widetilde{\varphi}\circ\widetilde{\varphi^{\prime}}$.
  4. (4)
    For all integers $m\geq 1$, we have $\widetilde{[m]_{X}}=[m^{2n-1}]_{X}$.
Proof.
Part (1) is clear from comparing degrees, given that $\widetilde{\varphi}\circ\varphi=[d]_{X}$ and $\textnormal{deg}([d]_{X})=d^{2n}$. As for part (2), note that
$\left([d^{2n-2}]_{Y}\circ\varphi\right)\circ\widetilde{\varphi}$ $=[d^{2n-2}]_{Y}\circ[d]_{Y}$
$=[d^{2n-1}]_{Y}$
$=\widetilde{\widetilde{\varphi}}\circ\widetilde{\varphi}$.
The claim then follows from Lemma 2.5. Part (3) is seen as follows:
$\left(\widetilde{\varphi}\circ\widetilde{\varphi^{\prime}}\right)\circ\left(\varphi^{\prime}\circ\varphi\right)$ $=\widetilde{\varphi}\circ[d^{\prime}]_{Y}\circ\varphi$
$=[dd^{\prime}]_{X}$
$=\widetilde{\varphi^{\prime}\circ\varphi}\circ\left(\varphi^{\prime}\circ\varphi\right).$
Part (4) is a now routine verification from the above parts. ∎
2.2.2. Separability and biseparability
Let $\varphi:X\rightarrow Y$ be an isogeny of abelian varieties over $K$. We set
$\textnormal{deg}_{\textnormal{sep}}(\varphi):=[\overline{K}(X):\varphi^{*}(\overline{K}(Y))]_{\textnormal{sep}}$
and
$\textnormal{deg}_{\textnormal{insep}}(\varphi):=[\overline{K}(X):\varphi^{*}(\overline{K}(Y))]_{\textnormal{insep}}$
to be the separability and inseparability degrees, respectively, of the extension of function fields $\overline{K(X)}/\varphi^{*}(\overline{K(Y)}).$ Equivalently, $\textnormal{deg}_{\textnormal{sep}}(\varphi)$ is equal to the number of closed points of $\textnormal{ker}(\varphi)$ and
$\textnormal{deg}_{\textnormal{sep}}(\varphi)\cdot\textnormal{deg}_{\textnormal{insep}}(\varphi)=\textnormal{deg}(\varphi).$
We call $\varphi$ separable if the field extension $K(\varphi)/K$ is separable, i.e., if $\textnormal{deg}_{\textnormal{insep}}(\varphi)=1$, and purely inseparable if this field extension is purely inseparable, i.e., if $\textnormal{deg}_{\textnormal{sep}}(\varphi)=1$. We call $\varphi$ biseparable if both $\varphi$ and $\widetilde{\varphi}$ are separable. If $p=\textnormal{char}(K)=0$, then all isogenies of abelian varieties over $K$ are separable and hence also biseparable. On the other hand, we have
Proposition 2.8.
[Mum08, Proposition, p.64] The multiplication-by-$n$ isogeny $n]:X\rightarrow X$ on an abelian variety $X$ over a field $K$ of characteristic $p>0$ is of degree $n^{2\cdot\textnormal{dim}(X)}$, and is separable if and only if $n$ is coprime to $p$.
Corollary 2.9.
An isogeny $\varphi$ of degree $d$ over a field $K$ of characteristic $p>0$ is biseparable if and only if $d$ is coprime to $p$.
Proof.
This follows from Proposition 2.8, and the fact that $\widetilde{\varphi}\circ\varphi=[d]_{X}$. ∎
Remark 2.10.
Our definition of biseparable is a generalization of that from [GP22]. In an effort to be as general as possible and not require polarizations, we believe that working with $\widetilde{\varphi}$ rather than the dual isogeny $\varphi^{\vee}$ of $\varphi$ is the right choice in this work.
If $p=\textnormal{char}(K)>0$, then separability does not imply biseparability. As a primary example, for $X/K$ an abelian variety the Frobenius homomorphism
$F_{X}:X\rightarrow X^{(p)}$
is purely inseparable of degree $p^{\textnormal{dim}(X)}$ [EvdGM, Proposition 5.15]. The Verschiebung homomorphism
$V_{X}:X^{(p)}\rightarrow X$
is also an isogeny of degree $p^{\textnormal{dim}(X)}$. We have that $V_{X}\circ F_{X}=[p]_{X}$ and $F_{X}\circ V_{X}=[p]_{X^{(p)}}$, and these homomorphisms are Cartier dual to one-another [EvdGM, Proposition 5.19, Proposition 5.20]. The separability degree of $V_{X}$ is intimately related to the structure of the $p$-power torsion of $X$.
Proposition 2.11.
Suppose $p=\textnormal{char}(K)>0$, and let $X/K$ be an abelian variety. There then exists an integer $r=r(X)$ with $0\leq r\leq\textnormal{dim}(X)$ such that
$X[p^{m}]\simeq\left(\mathbb{Z}/p^{m}\mathbb{Z}\right)^{r}$
for all $m\geq 0$. Explicitly, we have that $p^{r}=\textnormal{deg}_{\textnormal{sep}}(V_{X})$. Moreover, this quantity is invariant under isogeny; if $X$ is isogeneous to $Y$ then $r(X)=r(Y)$.
Proof.
This is [EvdGM, Proposition 5.22]. The fact that $p^{r}=\textnormal{deg}_{\textnormal{sep}}(V_{X})$ is not explicitly stated in their proposition, but is inherent in their proof. ∎
It is a result of Deuring (see [Mum08, p. 217]) that for an elliptic curve $E/K$, non-isotriviality is equivalent to $r(E)=1$. Both properties are therefore equivalent to $V_{E}$ being separable. In this context, $F_{E}$ and $V_{E}$ are dual isogenies, and $V_{E}$ is a separable isogeny which is not biseparable.
Suppose that $E_{1},\ldots,E_{n}$ are non-isotrivial, pairwise non-isogeneous elliptic curves over $K$ and that $p=\textnormal{char}(K)>0$. We then have
$\left(E_{1}\times\cdots\times E_{n}\right)[p^{m}]\simeq E_{1}[p^{m}]\times\cdots\times E_{n}[p^{m}]\simeq\left(\mathbb{Z}/p^{m}\mathbb{Z}\right)^{n}.$
Moreover, if $A$ is an abelian variety isogeneous to the product $E_{1}\times\ldots\times E_{n}$, then by Proposition 2.11 we have $r(A)=n=\textnormal{dim}(A)$. In this setting, therefore, the Verschiebung isogeny $V_{A}$ is separable. This will be the context of the results in this work.
Note that, in the above context, $A$ cannot be defined over a finite field; if there were some finite extension of $k$ over which we had a model for $A$, then there would exist some further finite extension over which an isogeny $\varphi:A\rightarrow E_{1}\times\cdots\times E_{n}$ is defined. The product $E_{1}\times\cdots\times E_{n}$ must then be defined over said extension, and upon a further finite extension each $E_{i}$ would be defined, contradicting our non-isotriviality assumption.
Remark 2.12.
The preceding paragraph addresses the case in which $A$ is isogeneous to a totally decomposable abelian variety. For simple abelian varieties over a field $K$ of positive characteristic, a result due to Tate and Grothendieck states that a simple abelian variety $X/K$ has a model over a finite field if and only if it is of CM type [Mum08, p. 220]. For elliptic curves, “CM type” is equivalent to isotriviality.
Lemma 2.13.
Let $X$ and $Y$ be abelian varieties defined over a finite extension $L$ of $K$ which have no abelian subvarieties of CM type. Let $\varphi:X\rightarrow Y$ be a biseparable isogeny. Then letting $H=\textnormal{ker}(\varphi)(\overline{K})$, we have that $L(H)/L$ is a separable field extension.
Proof.
This proof follows similarly to that in [GP22, Lemma 4.6]: we know that $H\subseteq X[d]$, where $d=\textnormal{deg}(\varphi)$, and because $d$ is coprime to $p$ it follows from [ST68, §1] that $L(X[d])/L$ is separable. Hence, $L(H)/L$ is separable. ∎
3. An isogeny degree bound
The following statement, providing an upper bound on the order of a $G_{K}$-stable subgroup of an elliptic curve over $K$ in terms of just the genus of $K$, is [GP22, Proposition 6.5]. This comes from a clever observation that, in the function field setting, a $K$-rational point on the modular curve $X_{0}(d)$ induces a morphism $C\to X_{0}(d)$ from our smooth model $C$ of $K$ to $X_{0}(d)$ over the constant field $k$.
Lemma 3.1.
Let $E$ be a non-isotrivial elliptic curve over $K$. Let $H\subseteq E$ be a cyclic, $G_{K}$-stable subgroup of order $d$. If $d$ is coprime to $\textnormal{char}(K)$, then
$d\leq 49\cdot\textnormal{max}\{1,g(C)\}.$
In their work, Griffon–Pazuki use Lemma 3.1 to obtain the following degree bounds for biseparable isogenies of elliptic curves over function fields.
Proposition 3.2.
[GP22, Proposition 6.7] Let $E_{1},E_{2}$ be two non-isotrivial elliptic curves over a function field $K$. Assume that there exists a biseparable isogeny $\varphi:E_{1}\rightarrow E_{2}$ over $K$. Then there exists a biseparable isogeny $\varphi_{0}:E_{1}\rightarrow E_{2}$ over $K$ with
$\deg{\varphi_{0}}\leq 49\cdot\max\{1,g(K)\}.$
Proposition 3.3.
Let $K$ be a function field and let $L/K$ be a finite extension. Let $E_{1},E_{2}$ be two non-isotrivial elliptic curves over $L$, and assume that there exists a biseparable isogeny $\varphi:E_{1}\rightarrow E_{2}$ over $L$. Then there exists a biseparable isogeny $\varphi_{0}:E_{1}\rightarrow E_{2}$ with
$\deg{\varphi_{0}}\leq[L:K]\cdot\min\{\mathrm{h}_{\textnormal{mod}}(E_{1}),\mathrm{h}_{\textnormal{mod}}(E_{2})\}.$
Proof.
This follows from the same argument given for [GP22, Proposition 6.7], but using the cruder degree bound of [GP22, Proposition 6.6] in place of [GP22, Proposition 6.5]. ∎
We leverage Lemma 3.1 and the argument in the proof of Proposition 3.2 to obtain analogous results for a product of two elliptic curves.
Proposition 3.4.
Let $E_{1},E_{2}$ be non-isotrivial, non-isogenous elliptic curves over $K$, and let $X$ be an abelian surface over $K$. Assume that there exists a biseparable isogeny $\varphi:E_{1}\times E_{2}\rightarrow X$ over $K$. Then there exists a biseparable isogeny $\varphi_{0}:E_{1}\times E_{2}\rightarrow X$ over $K$ with
$\textnormal{deg}(\varphi_{0})\leq 9\cdot 49^{2}\cdot\textnormal{max}\{1,g(K)^{2}\}.$
Proof.
The structure of our proof will be similar to that of [GP22, Proposition 6.7], with the added task of relating our isogeny degree bound to induced points on modular curves. We set
$d_{0}:=\textnormal{min}\{d\in\mathbb{Z}^{+}\mid\exists\textnormal{ a biseparable isogeny }E_{1}\times E_{2}\rightarrow X\text{ of degree }d\},$
which is well-defined by our hypothesis that there exists a biseparable isogeny between these abelian surfaces. Let $\varphi_{0}:E_{1}\times E_{2}\rightarrow X$ be a choice of biseparable isogeny realizing this degree. By Corollary 2.9, we have $p\nmid d_{0}$.
Letting $0_{1}$ and $0_{2}$ denote, respectively, the identity elements of $E_{1}$ and $E_{2}$, let
$E_{1}^{\prime}$ $:=\varphi_{0}\left(E_{1}\times\{0_{2}\}\right)\quad\textnormal{ and }$
$E_{2}^{\prime}$ $:=\varphi_{0}\left(\{0_{1}\}\times E_{2}\right).$
Each elliptic curve $E_{i}^{\prime}\leq X$ is an abelian subvariety of $X$, and we have isogenies
$\varphi_{1}:$ $E_{1}\longrightarrow E_{1}^{\prime}$
$P\longmapsto\varphi_{0}(P,0_{2}),\textnormal{ and }$
$\varphi_{2}:$ $E_{2}\longrightarrow E_{2}^{\prime}$
$Q\longmapsto\varphi_{0}(0_{1},Q).$
Note that $\textnormal{ker}(\varphi_{1})(\overline{K})\times\{0_{2}\}\leq\textnormal{ker}(\varphi_{0})(\overline{K})$ and $\{0_{1}\}\times\textnormal{ker}(\varphi_{2})(\overline{K})\leq\textnormal{ker}(\varphi_{0})(\overline{K})$. Moreover, if $(P,Q)\in\textnormal{ker}(\varphi_{0})(\overline{K})$, then with $0_{X}$ denoting the identity of $X$ we find
$\varphi_{1}(P)+\varphi_{2}(Q)=\varphi_{0}(P,Q)=0_{X},$
or equivalently $\varphi_{1}(P)=\varphi_{2}(-Q)$. Consequently, we have $\varphi_{1}(P)\in E_{1}^{\prime}\cap E_{2}^{\prime}$. For a given point $R\in E_{1}^{\prime}\cap E_{2}^{\prime}$, there are $d_{i}$ points $P\in E_{i}$ with $\varphi_{i}(P)=R$, and each pair of points $P\in E_{1}$ and $Q\in E_{2}$ with $\varphi_{1}(P)=\varphi_{2}(Q)=R$ provides a distinct point $(P,Q)\in\textnormal{ker}(\varphi_{0})(\overline{K})$. The fact that $E_{1}$ and $E_{2}$ are non-isogenous provides that $|E_{1}^{\prime}\cap E_{2}^{\prime}|\leq 9$, which leads us to
(3.1) $d_{0}=|\textnormal{ker}(\varphi_{0})(\overline{K})|=d_{1}d_{2}\cdot|E_{1}^{\prime}\cap E_{2}^{\prime}|\leq 9\cdot d_{1}d_{2}.$
We claim that both $\varphi_{1}$ and $\varphi_{2}$ are cyclic isogenies of elliptic curves. Suppose, for example, that $\varphi_{1}$ is not cyclic, so there exists some $m\mid d_{1}$ with $m>1$ such that $\varphi_{1}$ factors through the endomorphism $m]_{E_{1}}:E_{1}\to E_{1}$. It follows that $\varphi_{0}$ factors through the endomorphism
$\eta:=[m]_{E_{1}}\times\textnormal{Id}_{E_{2}}\in\operatorname{End}(E_{1}\times E_{2})\cong\mathbb{Z}\times\mathbb{Z},$
and hence by Proposition 2.4 there exists an isogeny $\psi:E_{1}\times E_{2}\rightarrow X$ with $\varphi_{0}=\psi\circ\eta$. By comparing degrees, we find that $\psi$ is both biseparable and of degree strictly less than $d_{0}$, contradicting the minimality of the degree of $\varphi_{0}$.
Now, by the same arguments as in the proof of [GP22, Proposition 6.7] and the fact that $d_{1},d_{2}\mid d_{0}$, both $\varphi_{1}$ and $\varphi_{2}$ are cyclic and biseparable with kernel stable under the action of $G_{K}$. Thus $\varphi_{i}$ induces a $K$-rational point on $X_{0}(d_{i})$ for $i=1,2$. Lemma 3.1 then provides
$d_{1},d_{2}\leq 49\cdot\textnormal{max}\{1,g(K)\},$
and from Equation 3.1 we reach the stated result. ∎
Proposition 3.5.
Let $K$ be a function field and $L/K$ a finite extension. Let $E_{1},E_{2}$ be non-isotrivial, non-isogenous elliptic curves over $L$, and let $X$ be an abelian surface over $L$. Assume that there exists a biseparable isogeny $\varphi:E_{1}\times E_{2}\rightarrow X$ over $L$. Then there exists a biseparable isogeny $\varphi_{0}:E_{1}\times E_{2}\rightarrow X$ over $L$ with
$\textnormal{deg}(\varphi_{0})\leq 9\cdot[L:K]^{2}\cdot\mathrm{h}_{\textnormal{mod}}(E_{1})\cdot\mathrm{h}_{\textnormal{mod}}(E_{2}).$
Proof.
One need only apply [GP22, Proposition 6.6] in place of Lemma 3.1 in the proof of Proposition 3.4, as was done to obtain Proposition 3.3 compared to Proposition 3.2 in the elliptic curve case. ∎
While Proposition 3.2 and Proposition 3.4 have the pleasing feature of uniformity, depending only on the genus of $K$, there are situations for which the non-uniform versions Proposition 3.3 and Proposition 3.5 are better tuned. For instance, Griffon–Pazuki use Proposition 3.3 to prove a bound on sizes of isogeny classes over fixed degrees [GP22, Proposition 6.12], for which Proposition 3.2 is too crude as genera may be unbounded over a family of extensions of $K$ of a fixed degree.
4. Effective surjectivity for products
In this section, we apply the results of Section 3 to obtain an effective open image-type result for Galois representations of products of elliptic curves over a function field $K$.
An effective result in the case of a single elliptic curve was obtained by Cojocaru–Hall [CH05, Theorem 1.1], and we will make use of their theorem here. To state it, we recall the following setup. For $E$ an elliptic curve over $K$ and $\ell$ a rational prime not equal to $\textnormal{char}(K)$, let
$\rho_{E,\ell}:G_{K}\rightarrow\operatorname{GL}(E[n])\simeq\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$
denote the mod-$\ell$ Galois representation of $E$, obtained via the action of $G_{K}$ on the $\ell$-torsion points of $E$. (The isomorphism with $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ here is non-canonical, depending on a choice of basis for $E[\ell]$.) Letting $\chi_{\ell}:G_{K}\to(\mathbb{Z}/\ell\mathbb{Z})^{\times}$ denote the $\ell^{\textnormal{th}}$ cyclotomic character of $K$, we have that $\textnormal{det}\circ\rho_{E,\ell}=\chi_{\ell}$. Hence, the image $G_{\ell}$ of $\rho_{E,\ell}$ must be a subgroup of the group $\Gamma_{\ell}\leq\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ determined by the short exact sequence
$0\rightarrow\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})\rightarrow\Gamma_{\ell}\rightarrow\chi_{\ell}(G_{K})\rightarrow 0.$
That is, $\Gamma_{\ell}$ is the inverse image in $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ of the determinant map. Following [CH05, §1], we call
$H_{\ell}:=G_{\ell}\cap\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})\subseteq\Gamma_{\ell}\cap\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})=\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$
the geometric Galois group of $K(E[\ell])/K$. The fixed field of $H_{\ell}$ corresponds to the constant extension $(K(E[\ell])\cap\overline{k})K/K$ obtained by adjoining a primitive $\ell^{\textnormal{th}}$ root of unity, and is invariant under a change in $K$ by a constant extension in this setup (hence the adjective “geometric” here). In other words, we have $G_{\ell}=H_{\ell}$ upon a base change of $C$ from $k$ to a cyclotomic extension of $k$, and so we restrict our attention to $H_{\ell}$ for a geometric result. Specifically, the effective result of Cojocaru–Hall states that $H_{\ell}$ is all of $\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ for sufficiently large $\ell$, where “sufficiently large” depends only on the genus of $K$.
Theorem 4.1.
[CH05, Theorem 1.1] Let $E$ be a non-isotrivial elliptic curve over a function field $K$ as above of genus $g$. Let
$e_{2}(q):=\begin{cases}1\quad&\text{ if }q\equiv 1\pmod{4}\\ -1\quad&\text{ otherwise,}\end{cases}\quad\quad e_{3}(q):=\begin{cases}1\quad&\text{ if }q\equiv 1\pmod{3}\\ -1\quad&\text{ otherwise,}\end{cases}\quad$
and set
$c(g):=2+\max\left\{q\textnormal{ prime }\bigm|\dfrac{q-(6+3e_{2}(q)+4e_{3}(q))}{12}\leq g\right\}.$
We then have for all primes $\ell\geq c(g)$ with $\ell\neq\textnormal{char}(K)$ that $H_{\ell}=\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$.
To get an effective surjectivity bound for products, we follow the techniques of Masser–Wüstholz [MW93a]. In the referenced work, the authors use an upgraded version [MW93a, Lemma 2.2] of an isogeny degree bound from prior work of theirs [MW93b] both to prove a result for a single elliptic curve over a number field [MW93a, Theorem], analogous to Theorem 4.1, and to bootstrap this version to products of an arbitrary number of elliptic curves [MW93a, Proposition 1]. We will do the same given our isogeny degree bound in the function field setting Proposition 3.5. The use of this degree bound, rather than Proposition 3.4, yields a result which is not uniform in the genus of $K$. We address why we use the weaker bound for our proof in Remark 4.4.
The following algebraic result of Masser–Wüstholz, which is a slight generalization of a result of Serre [Ser72, Lemme 8], will be useful in the course of our proof.
Lemma 4.2.
[MW93a, Lemma 5.1] For a prime $\ell\geq 5$, let $e$ be a divisor of $\ell-1$, let $V,V^{\prime}$ be vector spaces of dimension $2$ over $\mathbb{F}_{\ell}$, and let $B,B^{\prime}$ be the subgroups of $\operatorname{GL}(V)$ and $\operatorname{GL}(V^{\prime})$, respectively, consisting of all elements whose determinants are $e^{\textnormal{th}}$ powers. Let $D$ be the subgroup of $B\times B^{\prime}$ consisting of all $(b,b^{\prime})$ with $\det{b}=\det{b^{\prime}}$, and suppose $H$ is a subgroup of $D$ in $B\times B^{\prime}$ whose projections to each factor are surjective. Then if $H\neq D$ there is an isomorphism $f$ from $V$ to $V^{\prime}$ and a character $\chi$ of $H$ with $\chi^{2}=1$ such that
$b^{\prime}=\chi(h)fbf^{-1}$
for all $h=(b,b^{\prime})$ in $H$.
To state our theorem, we recall the following setup from Section 1. Let $E_{1},\ldots,E_{n}$ with $n\geq 2$ be non-isotrivial, pairwise non-isogenous elliptic curves over a function field $K$ of genus $g$, and let
$A=E_{1}\times\cdots\times E_{n}.$
For $\ell$ a prime, let
$\rho_{A,\ell}:G_{K}$ $\longrightarrow\operatorname{GL}(E_{1}[\ell])\times\cdots\times\operatorname{GL}(E_{n}[\ell])\simeq\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}$
$\sigma$ $\longmapsto(\rho_{E_{1},\ell}(\sigma),\ldots,\rho_{E_{n},\ell}(\sigma)),$
and let $\Psi_{\ell}:=\textnormal{Image}(\rho_{A,\ell})\cap\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}$.
Theorem 4.3.
Assume the setup of the prior paragraph, and set
$C(g)=3176523\cdot\max\{1,g(K)^{3/2}\}.$
For all primes $\ell$ not equal to $\textnormal{char}(K)$ satisfying $\ell>C(g)$, we have $\Psi_{\ell}=\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}$.
Proof.
We closely follow the proof structure of [MW93a, Proposition 1]. The dependence on the fact that the authors are working over a number field in the proof of this referenced result lies only in the fact that they use their prior results specific to that setting, which we will replace with Theorem 4.1 and Proposition 3.5.
First, we handle the $n=2$ case of a product of two elliptic curves. Let $E$ and $E^{\prime}$ be two non-isotrivial, non-isogenous elliptic curves over $K$, and assume that $\ell\geq c(g)$. Note that it follows that $\ell\geq c(0)=17$. Let $H=\textnormal{Image}(\rho_{A,\ell})$ and let $V=E[\ell]$ and $V^{\prime}=E^{\prime}[\ell]$. By Theorem 4.1 and our choice of $\ell$, we know that $H$ projects surjectively onto $\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ in each factor. We can then apply Lemma 4.2 with our choice of $H,V,$ and $V^{\prime}$ and with $B,B^{\prime}\simeq\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$.
If $H=D$ then we are done. Otherwise, our application of Lemma 4.2 tells us that we have an isomorphism $f:V\to V^{\prime}$ and a character $\chi_{0}$ on $G_{K}$ (induced from the character $\chi$ on $H$ given by the lemma) such that, for every $\sigma\in G_{K}$,
$\rho_{E^{\prime},\ell}(\sigma)=\chi_{0}(\sigma)f\rho_{E,\ell}(\sigma)f^{-1}.$
This means that $f$ induces a $G_{K}$-module isomorphism, which we will denote by $f_{\chi_{0}}$, from $E[\ell]$ to $E^{\prime}_{\chi_{0}}[\ell]$, where $E^{\prime}_{\chi_{0}}$ is the quadratic twist of $E^{\prime}$ by $\chi_{0}$; for every $\sigma\in G_{K}$
$\rho_{E^{\prime}_{\chi_{0}},\ell}(\sigma)=f_{\chi_{0}}\rho_{E,\ell}(\sigma)f_{\chi_{0}}^{-1}.$
Letting
$\Gamma:=\{(x,f_{\chi_{0}}x)\in E\times E^{\prime}_{\chi_{0}}\mid x\in E[\ell]\},$
we note that $\Gamma$ is a subgroup of $E\times E^{\prime}_{\chi_{0}}$ which is stable under the action of $G_{K}$. Indeed, if $\sigma\in G_{K}$ and we set $y=\rho_{E,\ell}(\sigma)x$, then we compute
$\sigma(x,f_{\chi_{0}}x)=(\rho_{E,\ell}(\sigma)x,\rho_{E^{\prime}_{\chi_{0}},\ell}(\sigma)f_{\chi_{0}}x)=(y,f_{\chi_{0}}y)\in\Gamma.$
Therefore, the abelian surfaces $A=E\times E^{\prime}_{\chi_{0}}$ and $X=(E\times E^{\prime}_{\chi_{0}})/\Gamma$ are both defined over $K$ and are isogenous via the natural quotient map $\pi:A\to X$.
By our assumptions on $E$ and $E^{\prime}$, and the fact that $\textnormal{deg}(\pi)=\ell^{2}$ is relatively prime to $\textnormal{char}(K)$ by our restriction on $\ell$, Proposition 3.4 provides the existence of an isogeny $\psi_{0}:A\to X$ with $\textnormal{deg}(\psi_{0})\leq 21609\cdot\max\{1,g(K)^{2}\}$. We then have an endomorphism
$\epsilon=\widetilde{\psi_{0}}\circ\pi\in\textnormal{End}_{\overline{K}}(A)\simeq\mathbb{Z}^{2},$
of degree $\ell^{2}\cdot\textnormal{deg}(\psi_{0})^{3}$, which we can represent by matrix $\begin{pmatrix}a&0\\ 0&b\end{pmatrix}$ with $a,b\in\mathbb{Z}$. As $\pi$ annihilates $\Gamma$ so too does $\epsilon$, and so for all $x\in E[\ell]$ we have
$ax=bf_{\chi_{0}}x=0.$
This means that the multiplication by $a$ map on $E$ and the multiplication by $b$ map on $E^{\prime}_{\chi_{0}}$ both annihilate $\ell$-torsion (as $f_{\chi_{0}}$ is an isomorphism), and so $\ell\mid a$ and $\ell\mid b$. Moreover, $\ell^{4}\mid a^{2}b^{2}=\textnormal{deg}(\epsilon)$, so $\ell^{2}\mid\textnormal{deg}(\psi_{0})^{3}$. We can then rule this case out by assuming that
$\ell>3176523\cdot\max\{1,g(K)^{3/2}\}.$
Thus, for all primes
$\ell$ $>\max\left\{c(g),3176523\cdot\max\{1,g(K)^{3/2}\}\right\}$
$=3176523\cdot\max\{1,g(K)^{3/2}\}$
we have the desired equality $\Psi_{\ell}=\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})\times\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$.
We now obtain the general case from the $n=2$ case as follows. Assume the inequality $\ell>C(g)$. The image of the map from $\Psi_{\ell}$ to the product of the form $\operatorname{SL}(E_{i})\times\operatorname{SL}(E_{j})$ for any $1\leq i<j\leq n$, via projection onto the two corresponding factors of $\Psi_{\ell}\leq\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}$, is then surjective by the $n=2$ case. As in the proof of [MW93a, Proposition 1], we now use the result of Ribet [Rib76, Lemma 5.2.2] that implies that $\Psi_{\ell}=\operatorname{SL}_{2}(\mathbb{Z}/n\mathbb{Z})^{n}$ (as $\ell>5$). ∎
Remark 4.4.
We thank Ariel Weiss for pointing out to us that the character $\chi_{0}$ in the proof of Theorem 4.3 yields an isogeny over $K$ with source abelian surface $E\times E^{\prime}_{\chi}$. Serre notes that the isomorphism $f$ yields an isomorphism between the $\ell$-torsion subgroups of $E$ and the twist $E^{\prime}_{\chi_{0}}$ [Ser72, p. 328], but instead of using the associated isogeny of abelian surfaces over $K$ as in our proof Masser–Wüstholz trivialize $\chi_{0}$ to get an isogeny over the at most quadratic extension $K^{\chi_{0}}/K$. Using their isogeny estimates, this line of argument only costs them a factor of $2$ in their final bound compared to the one we use, as it would for our version Theorem 4.5 based on our bound from Proposition 3.5. For us, the argument over $K$ allows us to forego other possible arguments limiting the ramification of the characters that show up, which would be needed for the uniform version Theorem 4.3; it is not immediately clear that the quadratic extensions $K^{\chi_{0}}/K$ obtained as one ranges over $\ell$ have uniformly bounded genus.
We next state a version of our effective surjectivity result in terms of the modular heights of the elliptic curve factors. This could be useful in the sense that the bound $C^{\prime}(g,E_{1},\ldots,E_{n})$ defined below can be smaller than the bound $C(g)$ from Theorem 4.3 for certain individual product abelian varieties.
Theorem 4.5.
Assume the above setup, and set
$C^{\prime}=C^{\prime}(g,E_{1},\ldots,E_{n})=\max\left\{c(g),27\cdot\max_{1\leq i<j\leq n}\left\{\mathrm{h}_{\textnormal{mod}}(E_{i})\cdot\mathrm{h}_{\textnormal{mod}}(E_{j})\right\}^{3/2}\right\}.$
For all primes $\ell$ not equal to $\textnormal{char}(K)$ satisfying $\ell>C^{\prime}$, we have $\Psi_{\ell}=\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}$.
Proof.
The only change necessary compared to the proof of Theorem 4.3 is the use of Proposition 3.5 in place of Proposition 3.4. ∎
Remark 4.6.
It is also tempting to say that Theorem 4.5 would be useful in place of Theorem 4.3 in cases where one works with a family of products of elliptic curves of bounded modular height over a family of function fields of unbounded genus, but this version still has dependence on the genus coming from the appearance of the constant $c(g)$ from the result of Cojocaru–Hall. If one were to replace the Cojocaru–Hall result with an explicit surjectivity result for elliptic curves which only depends on the modular height, though, then this would indeed be the case.
5. An average result over families of products
In this section, we prove results on exceptional primes for specializations of families of products of elliptic curves over $\mathbb{Q}$. Throughout, we let $n\geq 2$ be an integer and let $A_{1}(t),B_{1}(t),A_{2}(t),B_{2}(t),\ldots,A_{n}(t),B_{n}(t)$ be polynomials in $\mathbb{Z}[t]$ such that the elliptic curve
$E_{i}:y^{2}=x^{3}+A_{i}(t)x+B_{i}(t)$
over $\mathbb{Q}(t)$ is non-isotrivial (i.e., has $j$-invariant $j(E_{i})\not\in\overline{\mathbb{Q}}$) for all $i\in\{1,2,\ldots,n\}$, and such that $E_{i}$ and $E_{j}$ are not isogenous over $\overline{\mathbb{Q}(t)}$ for all $1\leq i<j\leq n$. We let
$A=E_{1}\times E_{2}\times\cdots\times E_{n}$
be the product abelian surface over $\mathbb{Q}(t)$, and for $t_{0}\in\mathbb{Q}$ we let $E_{i,t_{0}}$ and $A_{t_{0}}$ denote the corresponding specializations of $E_{i}$ and of $A$, respectively. Note that our assumptions imply that the geometric endomorphism ring of $A$ is $\textnormal{End}_{\overline{K}}(A)\simeq\mathbb{Z}^{n}$, and that for a given $t_{0}\in\mathbb{Q}$ the same holds for $A_{t_{0}}$ so long as
  1. (1)
    $A_{t_{0}}$ is an abelian variety of dimension $n$, i.e., the specialization of $\Delta_{i}(t):=-16[4A_{i}(t)^{3}+27B_{i}(t)^{2}]$ to $t_{0}$ is nonzero for all $i\in\{1,2,\ldots,n\}$, and
  2. (2)
    $E_{i,t_{0}}$ and $E_{j,t_{0}}$ are not isogenous over $\overline{\mathbb{Q}}$ for all $1\leq i<j\leq n$ and each $E_{i,t_{0}}$ is non-CM.
Condition (1) is satisfied for all but finitely many specializations, and condition (2) is also satisfied by most specializations. For the latter claim, see the result [Mas96, Theorem] on endomorphisms under specialization as well as Masser’s discussion in this work [Mas96, p. 461-462] referencing a result of Serre that condition (2) is satisfied for all rational specializations outside of a thin set. Let $S=S(E_{1},E_{2},\ldots,E_{n})$ denote the set of $t_{0}\in\mathbb{Q}$ for which at least one of Condition (1) and Condition (2) is not satisfied.
We first prove a result bounding specializations with prescribed Frobenius traces. We then use this as a stepping stone to get from Theorem 4.3 to Theorem 5.5. In the course of this proof, we will rely on Murty–Scherk’s effective version of the Chebotarev density theorem over function fields as recalled below:
Proposition 5.1.
[KMS94, Theorem 2] Let $X\rightarrow Y$ be a finite Galois cover of smooth projective curves over $K=\mathbb{F}_{q}(t)$ with corresponding Galois group $G$. Let $m$ be such that the algebraic closure of $K$ in $K(X)$ is $\mathbb{F}_{q^{m}}$. Let $Z$ be the geometric ramification locus of this map, and let $\pi(r)$ denote the number of unramified closed points of $Y$ of degree $r$. Suppose that $\mathcal{C}^{q^{r}}\subseteq G$ is a conjugacy class whose restriction to $\mathbb{F}_{q^{m}}$ is $\textnormal{Frob}_{q^{r}}$, and let
$\pi_{\mathcal{C}^{q^{r}}}(r)=\#\{y\in Y\otimes_{K}\mathbb{F}_{q^{r}}\mid y\textnormal{ is unramified, }\textnormal{deg}(y)=r,\textnormal{Frob}_{y}\subseteq\mathcal{C}^{p}\}.$
Then
$\pi_{C^{q^{r}}}(r)=m\frac{|\mathcal{C}^{q^{r}}|}{|G|}\frac{q^{r/2}}{r}+O_{K,Z}\left(|{\mathcal{C}^{q^{r}}}|^{1/2}q^{r/2}\right).$
More specifically, we will use a special case of their theorem suitable to our needs and stated in terms of an extension of function fields rather than in terms of a covering of the corresponding curves (like that given in [CH05, Proposition 3.1], germane to the case of a single elliptic curve).
Corollary 5.2.
Let $K=\mathbb{F}_{p}(t)$, let $F/K$ be a tamely ramified, finite Galois extension, and let $U$ be the open complement of the set $Z$ of ramified places of $K$. Let $G=\operatorname{Gal}(F/K)$ and suppose that $\mathcal{C}\subseteq G$ is a union of conjugacy classes. Let $G^{p}:=\{p\in G\mid g|_{\mathbb{F}_{p}}=\textnormal{Frob}_{p}\}$ and let $\mathcal{C}^{p}=G^{p}\cap\mathcal{C}$. Then
$\{\nu\in\mathcal{U}(\mathbb{F}_{p})\mid\textnormal{Fr}_{\nu}\subseteq\mathcal{C}^{p}\}=\dfrac{|\mathcal{C}^{p}|}{|G^{p}|}|\mathcal{U}(\mathbb{F}_{p})|+O_{K,Z}\left(|\mathcal{C}^{p}|^{1/2}p^{1/2}\right).$
Proof.
We apply Proposition 5.1 to the extension $F/K$ in the case $q^{r}=p$, using the fact that $\mathcal{C}^{p}$ is by construction a union of conjugacy classes. The only observation needed is that
$m|G^{p}|=|G|,$
where in our translation $m$ is such that $\mathbb{F}_{p^{m}}=F\cap\overline{\mathbb{F}_{p}}$. Indeed, let $\textnormal{res}:G\rightarrow\operatorname{Gal}\left(\mathbb{F}_{p^{m}}/\mathbb{F}_{p}\right)$ be the map restricting an automorphism to $\mathbb{F}_{p^{m}}$. Letting, for each integer $0\leq k\leq m-1$,
$G^{p^{k}}:=\{g\in G\mid\textnormal{res}(g)=\textnormal{Frob}_{p^{k}}\},$
we have $G^{1}=\textnormal{ker}(\textnormal{res})$ and $G^{1}$ is in bijection with $G^{p^{k}}$ via $\textnormal{Frob}_{p^{k}}$. Hence, $|G|=m\cdot|G^{1}|=m\cdot|G^{p}|$. ∎
For our theorem statement, we let $a_{p}(E)$ denote the Frobenius trace of an elliptic curve $E/\mathbb{Q}$, so $a_{p}(E)=p+1-|\overline{E}(\mathbb{F_{p}})|$. With $C(g)$ as in Theorem 4.3, note that
$C(0)=\max\{17,3176523\}=3176523.$
Theorem 5.3.
Let $E_{1},\ldots,E_{n}$ and $A$ as above. Let $p$ and $\ell$ be primes such that
  • $p\neq\ell$,
  • the specialization $\overline{E_{i}}$ of $E_{i}$ to $K=\mathbb{F}_{p}(t)$ is non-isotrivial for all $i\in\{1,\ldots,n\}$, and
  • $\ell\geq 3176533$.
For a fixed tuple $(\tau_{1},\ldots,\tau_{n})\in\mathbb{Z}^{n}$, let
$\mathcal{T}_{p}(\tau_{1},\ldots,\tau_{n}):=\#\left\{t_{0}\in\mathbb{F}_{p}\mid\Delta_{i}(t_{0})\not\equiv 0\pmod{p}\textnormal{ and }a_{p}(E_{i,t_{0}})\equiv\tau_{i}\pmod{\ell}\textnormal{ for }i\in\{1,\ldots,n\}\right\}.$
We then have
$\mathcal{T}_{p}(\tau_{1},\ldots,\tau_{n})=\frac{p}{\ell^{n}}+O_{A_{1},B_{1},\ldots,A_{n},B_{n}}\left(\ell^{n}p^{1/2}\right).$
Proof.
Our proof is similar to that of [CH05, Theorem 1.2] in the one-dimensional case. Let $S_{p}$ be the subset of places of $K$ at which the specialization $A_{p}$ of $A$ to $K$ has bad reduction, such that $S_{p}$ is finite of size bounded by a constant independent of $p$ and such that the extension $K(A_{p}[\ell])/K$ is unramified outside of $S_{p}$.
By hypothesis we have $\ell>C(0)$, so by Theorem 4.3 the geometric Galois group of $K(E_{1}[\ell],\ldots,E_{n}[\ell])/K$ is $\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}$. Thus, $G_{\ell}:=\textnormal{Gal}(K(E_{1}[\ell],\ldots,E_{n}[\ell])/K)$ is the unique subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})^{n}$ such that for each $(g_{1},\ldots,g_{n})\in G_{\ell}$ we have
$\det(g_{1})=\det(g_{2})=\ldots=\det(g_{n})$
and such that the image of the determinant on each factor of $G_{\ell}$ is $\langle p\rangle$.
In analogy to the notation used in the proof of [CH05, Theorem 1.2], set
$G_{\ell}^{p}$ $:=\{(g_{1},\ldots,g_{n})\in G_{\ell}\mid\det(g_{i})=p\},$
$\mathcal{C}_{\tau_{1},\ldots,\tau_{n}}$ $:=\{(g_{1},\ldots,g_{n})\in G_{\ell}\mid\textnormal{tr}(g_{i})=\tau_{i}\textnormal{ for }i\in\{1,\ldots,n\}\},\textnormal{ and}$
$\mathcal{C}_{\tau_{1},\ldots,\tau_{n}}^{p}$ $:=\mathcal{C}_{\tau_{1},\ldots,\tau_{n}}\cap G_{\ell}^{p}.$
We note that $\mathcal{C}_{\tau_{1},\ldots,\tau_{n}}^{p}$ is a union of conjugacy classes in $G_{\ell}$, namely those which are the products of the conjugacy classes for each factor considered in [CH05, Theorem 1.2].
We will now apply Corollary 5.2 to the extension $K(E_{1}[\ell],\ldots,E_{n}[\ell])/K$ with $\mathcal{C}^{p}=\mathcal{C}^{p}_{\tau_{1},\ldots,\tau_{n}}$. Note that in our application $\textnormal{deg}(Z)\leq\textnormal{deg}(S)$, so the implicit constant in the error term can be taken independent of $\ell$ and $p$. We have
$|G_{\ell}^{p}|=|\operatorname{SL}_{2}(\mathbb{Z}/\ell\mathbb{Z})|^{n}=\ell^{n}(\ell^{2}-1)^{n}.$
Because we have the relation $\mathcal{C}_{\tau_{1},\ldots,\tau_{n}}^{p}=\mathcal{C}_{\tau_{1}}^{p}\times\cdots\times\mathcal{C}_{\tau_{n}}^{p}$ where $\mathcal{C}_{\tau_{i}}^{p}$ are as defined in [CH05, Theorem 1.2], we may use their counts to obtain
$|\mathcal{C}_{\tau_{1},\ldots,\tau_{n}}^{p}|=\ell^{n}\prod_{i=1}^{n}(\ell+e_{i}\ell),$
where $e_{i}\in\{-1,0,1\}$ are as calculated in the referenced work of Cojocaru–Hall and only depend on $\delta_{i}=\tau_{i}^{2}-4p$ for each $i\in\{1,\ldots,n\}$. We thus obtain the stated result. ∎
The proof of Theorem 5.5, like the proof of [CH05, Theorem 1.3], will be an application of Gallagher’s $2$-dimensional large sieve. We recall this result below:
Lemma 5.4.
Let $\mathcal{A}$ be a subset of $\mathbb{Z}^{2}$, and $\mathcal{P}$ a set of rational primes. For each prime $p\in\mathcal{P}$, let $\Omega(p)$ be a subset of $\left(\mathbb{Z}/p\mathbb{Z}\right)^{2}$ with $\omega(p)$ elements. For $x>0$ and $\alpha=(\alpha_{1},\alpha_{2})\in\mathcal{A}$, let $\overline{\alpha}$ denote the pair $(\overline{\alpha_{1}},\overline{\alpha_{2}})\in\left(\mathbb{Z}/p\mathbb{Z}\right)^{2}$ of reductions modulo $p$ and set
$P(x)$ $:=\sum_{\begin{subarray}{c}p\in\mathcal{P}\\ p\leq x\end{subarray}}\dfrac{\omega(p)}{p^{2}}$
$P(\alpha,x)$ $:=\#\left\{p\leq x\mid p\in\mathcal{P}\textnormal{ and }\overline{\alpha}\in\Omega(p)\right\}.$
Then for any $z\geq x^{2}$,
$\sum_{\begin{subarray}{c}\alpha\in\mathcal{A}\\ |\alpha|\leq z|\end{subarray}}\left(P(\alpha,x)-P(x)\right)^{2}\ll z^{2}P(x),$
where $|\alpha|:=\max\{\alpha_{1},\alpha_{2}\}$.
For our theorem statement, we recall the following further setup beyond what we have already fixed: for $T>0$ a real number, set
$\mathcal{F}(T):=\left\{\frac{m}{n}\in\mathbb{Q}\setminus S\mid m,n\in\mathbb{Z},\gcd(m,n)=1,\textnormal{ and }\max\{|m|,|n|\}\leq T\right\}.$
For $\ell$ a rational prime set
$\mathcal{E}_{\ell}(T):=\left\{t_{0}\in\mathcal{F}(T)\mid\operatorname{Gal}(\mathbb{Q}(A_{t_{0}}[\ell])/\mathbb{Q})\subsetneq\Delta_{\ell}\right\}.$
Theorem 5.5.
Letting
$\mathcal{E}(T):=\bigcup_{\ell\geq 3176533}\mathcal{E}_{\ell}(T),$
we have
$\lim_{T\to\infty}\dfrac{|\mathcal{E}(T)|}{|\mathcal{F}(T)|}=0.$
Proof.
Let $\ell\geq 3176533$ be a fixed prime, and let $\delta,\tau_{1},\tau_{2},\ldots,\tau_{n}\in\mathbb{Z}/\ell\mathbb{Z}$ with $\delta\neq 0\pmod{\ell}$. For $t_{0}\in\mathcal{F}(T)$, let $N_{i,t_{0}}$ denote the conductor of $E_{i,t_{0}}$ and let
$\pi_{t_{0}}(x,\ell,\delta,\tau_{1},\ldots,\tau_{n})$
$:=\#\left\{p\leq x\mid p\nmid\ell N_{1,t_{0}}\cdots N_{n,t_{0}},p\equiv\delta\pmod{\ell},\textnormal{ and }a_{p}(E_{i,t_{0}})\equiv\tau_{i}\pmod{\ell}\textnormal{ for }i\in\{1,\ldots,n\}\right\}.$
Further let
$\mathcal{E}_{\ell,\delta,\tau_{1},\ldots,\tau_{n}}(T)$ $:=\left\{t_{0}\in\mathcal{F}(T)\mid\pi_{t_{0}}(x,\ell,\delta,\tau_{1},\ldots,\tau_{n})=0\textnormal{ for all }x>0\right\},$
$\mathcal{A}$ $:=\left\{(m,n)\in\mathbb{Z}^{2}\mid\gcd(m,n)=1\textnormal{ and }\frac{m}{n}\not\in S\right\},$
$\mathcal{P}$ $:=\left\{p\textnormal{ prime }\mid p\equiv\delta\pmod{\ell}\right\}$,
$\Omega(p)$ $:=\left\{(m,n)\in\left(\mathbb{Z}/p\mathbb{Z}\right)^{2}\mid\gcd(n,p)=1,p\nmid N_{1,m/n}\cdots N_{n,m/n},\textnormal{ and }a_{p}(E_{i,m/n})\equiv\tau_{i}\pmod{\ell}\right\}$.
One should think of $\mathcal{E}_{\ell,\delta,\tau_{1},\ldots,\tau_{n}}(T)$ as our set of exceptional specializations witnessed by determinant $\delta$ and Frobenius traces $\tau_{1},\ldots,\tau_{n}$ of $\mathcal{A}$ as our set of specializations with good reduction (including the endomorphism condition on the specialization), of $\mathcal{P}$ as those primes witnessing $\delta$ arising as the determinant of an element in the image of the Galois representation of the product, and of $\Omega(p)$ as those specializations which witness Frobenius traces $\tau_{1},\ldots,\tau_{n}$ modulo $p$. Applying Lemma 5.4 with the above yields
(5.1) $P\left(\sqrt{T}\right)^{2}\cdot|\mathcal{E}_{\ell,\delta,\tau_{1},\ldots,\tau_{n}}(T)|$ $=\sum_{\begin{subarray}{c}(m,n)\in\mathcal{A}\\ \max\{|m|,|n|\}\leq T\\ (m/n)\in\mathcal{E}_{\ell,\delta,\tau_{1},\ldots,\tau_{n}}(T)\end{subarray}}\left(\pi_{m,n}\left(\sqrt{T};\ell,\delta,\tau_{1},\ldots,\tau_{n}\right)-P\left(\sqrt{T}\right)\right)^{2}$
$\leq\sum_{\begin{subarray}{c}(m,n)\in\mathcal{A}\\ \max\{|m|,|n|\}\leq T\end{subarray}}\left(\pi_{m,n}\left(\sqrt{T};\ell,\delta,\tau_{1},\ldots,\tau_{n}\right)-P\left(\sqrt{T}\right)\right)^{2}$
$\ll T^{2}\cdot P\left(\sqrt{T}\right)\;.$
The first equality in Equation 5.1 follows from the definition of $\mathcal{E}_{\ell,\delta,\tau_{1},\ldots,\tau_{n}}(T)$, the second inequality follows from non-negativity of the summands, and the third is via Gallagher’s sieve.
We now seek a lower bound on $P\left(\sqrt{T}\right)$. Because $\ell>C(0)$, Theorem 5.3 provides
$\omega(p):=|\Omega(p)|=(p-1)\left(\frac{p}{\ell^{n}}+O_{A_{1},B_{1},A_{2},B_{2}}\left(\ell^{n}p^{1/2}\right)\right),$
and from here we compute
(5.2) $P\left(\sqrt{T}\right)$ $=\sum_{\begin{subarray}{c}p\leq\sqrt{T}\\ p\neq\ell\\ p\equiv\delta\pmod{\ell}\end{subarray}}\frac{\omega(p)}{p^{2}}$
$=\sum_{\begin{subarray}{c}p\leq\sqrt{T}\\ p\neq\ell\\ p\equiv\delta\pmod{\ell}\end{subarray}}\left(\frac{1}{\ell^{n}}+O_{A_{1},B_{1},A_{2},B_{2}}\left(\ell^{n}p^{-1/2}\right)\right)$
$=\frac{1}{\ell^{n}}\pi\left(\sqrt{T};\ell,\delta\right)+O_{A_{1},B_{1},A_{2},B_{2}}\left(\frac{T^{1/4}}{\log T}\right),$
with $\pi\left(x;\ell,\delta\right)$ denoting the number of primes $p\leq x$ with $p\equiv\delta\pmod{\ell}$.
We now proceed just as in the proof of [CH05, Theorem 1.3], but using the result of Masser–Wüstholz [MW93a, Proposition 1] for products of elliptic curves over number fields; we know that if $t_{0}\in\mathcal{E}_{\ell,\delta,\tau_{1},\ldots,\tau_{n}}(T)$ and thus $\ell$ is an exceptional prime for $A_{t_{0}}/\mathbb{Q}$, then there exist positive constants $c_{1},\gamma_{1}$ such that
$\ell\leq c_{1}\max\{\mathrm{h}_{\textnormal{mod}}(E_{1,t_{0}}),\mathrm{h}_{\textnormal{mod}}(E_{2,t_{0}})\}^{\gamma_{1}}.$
There then exists positive constants $c_{2}$ and $\gamma_{2}$ depending on $A_{1},B_{1},\ldots,A_{n},B_{n}$ such that
$\ell\leq c_{2}\left(\log T\right)^{\gamma_{2}}.$
From the Siegel–Walfisz theorem (see [IK04, Corollary 5.29]), we have that if $\ell\ll\left(\log{x}\right)^{B}$ for some real number $B>0$ then $\pi(x;\ell,\delta)\gg\frac{x}{\ell\log{x}}$. Thus, for $\ell$ and $t_{0}$ as above we have
$|\mathcal{E}_{\ell,\delta,\tau_{1},\ldots,\tau_{n}}(T)|=\frac{T^{2}}{P\left(\sqrt{T}\right)}\ll_{A_{1},B_{1},A_{2},B_{2}}T^{3/2}\log(T)^{\gamma_{3}}$
for some real number $\gamma_{3}>0$. We are now set to bound our full set of exceptional specializations of bounded height; we have
$|\mathcal{E}(T)|$ $\leq\sum_{\ell\geq 3176533}|\mathcal{E}_{\ell}(T)|$
$\leq\sum_{\ell\geq 3176533}\sum_{\begin{subarray}{c}\delta\in\left(\mathbb{Z}/\ell\mathbb{Z}\right)^{\times}\\ \tau_{1},\ldots,\tau_{n}\in\mathbb{Z}/\ell\mathbb{Z}\end{subarray}}|\mathcal{E}_{\ell,\delta,\tau_{1},\ldots,\tau_{n}}(T)|.$
The first sum is finite, and we know from above that $\mathcal{E}_{\ell,\delta,\tau_{1},\ldots,\tau_{n}}(T)$ is 0 if $\ell>c_{2}\left(\log{T}\right)^{\gamma_{3}}$. There then exists some positive real number $\gamma_{4}$ such that
$|\mathcal{E}(T)|\ll_{A_{1},B_{1},\ldots,A_{n},B_{n}}T^{3/2}\log(T)^{\gamma_{4}}.$
As $|\mathcal{F}(T)|\gg_{A_{1},B_{1},\ldots,A_{n},B_{n}}T^{2}$, our proof is complete. ∎